Three 15 Ohm resistors are connected in parallel and placed across a 30 V potential difference. a) What is the equivalent resistance of the parallel circuit? b) What is the current through the entire circuit? c) What is the current though each branch of the parallel circuit?


1
Expert's answer
2021-05-20T18:24:36-0400

a)


"\frac{1}{R}=\frac{1}{15}+\frac{1}{15}+\frac{1}{15}\\R=5\Omega"

b)


"I=\frac{V}{R'}=\frac{30}{15}=2\ A"